OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (sheraz12345):

find the slope of the line whose equation is given as the lines are paralell the slope would be same then find the equation using formula y-y1=m(x-x1)

OpenStudy (anonymous):

i have no idea how to do that

OpenStudy (sheraz12345):

do u know how to find slope from the given equation

OpenStudy (anonymous):

no

OpenStudy (sheraz12345):

Compare the given equation with y=mx+c by simplifying -3y=-12-4x divide by -3 on BS y=4+4/3x or \[y=\frac{ 4 }{ 3 }x+4\]

OpenStudy (sheraz12345):

hence my comparing with y=mx+c we get the slope BTW m is the slope so our slope=m=4/3

OpenStudy (sheraz12345):

Now we find the equation by using the formula \[y-y1=m(x-x1)\] insert the values \[y-4=\frac{ 4 }{ 3 }(x-(-3)\]

OpenStudy (sheraz12345):

by simplifying we get 3y -12=4x+12 3y-4x-24=0 Your Answer

OpenStudy (anonymous):

thses are the anwswers

OpenStudy (sheraz12345):

a part is the answer

OpenStudy (sheraz12345):

U got it @stupid12